OpenStudy (anonymous):

Q: Calculate the energy of a wavelength using this formua: E=h*f Key: E= energy H= Planck's constant (6.626*10^-34) units:js f= frequency Frequecny formula: 3.0*10^8 m/s (speed of light)/ wavelength (630nm) *NOTE: I've already done this, I just need someone to check my work. This is what I did: I found the frequency by: 3.0*10^8 m/s divided by 630nm. I got 476190.4762 for frequency (please help with units, I don't know how to put them together) Then I multiplied plancks constant with the frequency: 6.626*10^-34 * 476190.4762 to get: 15361905 (rounded) as the energy.

OpenStudy (anonymous):

COULD SOMEBODY PLEASE CHECK MY WORK, I have a lab report and I really want to get this right. :(

OpenStudy (phi):

you should use scientific notation when doing this problem 630 nm = 630 x 10^-9 m = 6.3 x 10^-7 m

OpenStudy (phi):

f= 3.0*10^8/6.3*10^-7 = 4.76*10^14 Hertz (1/sec)

OpenStudy (phi):

so your freq is too small

OpenStudy (anonymous):

So my error is not putting it in scientific notation?

OpenStudy (phi):

E= 6.626*10^-34 J-s * 4.76*10^14 1/sec = 3.15*10^-19 Joules

OpenStudy (phi):

I would not say error. Just very very bad way to do it. (too ugly)

OpenStudy (anonymous):

I won't say I'm not still confused, but thank you for showing me the "easier" way. :D

OpenStudy (phi):

I did not say it was easy. Just easier. First, you change 620 nm (nanometers) to meters. google for definition. you get as above. see if you can follow the arithmetic

OpenStudy (phi):

*630 nm